What is the ratio of 12 & 3/8 to 2 & 3/4?
Answer by stanbon(75887)

What is the ratio of 12 & 3/8 to 2 & 3/4?
-----------------
12 3/8 = (8*12+3)/8 = 99/8
------
2 3/4 = (4*2+3)/4 = 9/4
-------
Ans: (99/8)/(9/4) = 11/2 = 11:2
----------------------------------
Cheers,
Stan H.
====================